OpenStudy (anonymous):

The surface area S (in square meters) of a hot-air balloon is given by S(r) = 4π r2 where ris the radius of the balloon (in meters). If the radiusris increasing with time t(in seconds) according to the formula r(t) = (2/3)t^3 , t ≥ 0, find the surface area S of the balloon as a function of the time t.

OpenStudy (e.mccormick):

OK, how far have you gotten with this?

OpenStudy (anonymous):

nowhere, I have no clue what to do

OpenStudy (e.mccormick):

OK, had any function composition? Like \((f\circ g)(x)\) stuff?

OpenStudy (anonymous):

Nope, this is the first assignment from this chapter. I take math online so im pretty freaking confused.

OpenStudy (e.mccormick):

AAARGH! I had it mostly typed and then it booted me. Welp, time to try again.

OpenStudy (anonymous):

aw thank you!

OpenStudy (e.mccormick):

OK. I was going to go over composition really fast. Lets say I have \(f(x)=x-1\) and \(g(x)=(x+4)^2\). Then \((f\circ g)(x)\) means I plug the equation for g into the equation for f. It becomes this:\[f(g(x))=((x+4)^2)-1\]Your question is similar to this.

OpenStudy (e.mccormick):

Typed it in another window... just in case.

OpenStudy (anonymous):

Oh ok I vaguely remember doing that in high school

OpenStudy (e.mccormick):

Now you have: \(S(r) = 4\pi r^2\) and are told r is radius. Then they do a second formula for radius: \(r(t) = (2/3)t^3\) What they want is \((S\circ r)(t)\) which is also called \(S(r(t))\)

OpenStudy (anonymous):

would it look like 4pi ((2/3)t^3))^2 ?

OpenStudy (e.mccormick):

Yes, and you might do a little distribution. \[(x^m)^n=x^{m\times n}\]